The density of lead is 11g/cm 3 at 20 0 C. Find its density at 200 0 C


1
Expert's answer
2021-04-20T16:50:05-0400

The volume at "T_0 = 20\degree C" is:


"V_0 = \dfrac{m}{\rho_0}"

where "m" is the mass, and "\rho_0 = 11g/cm^3 = 11000\space kg/m^
3" is the density at "T_0".

The volume "V" at "T = 200\degree C" is given as follows (see https://en.wikipedia.org/wiki/Thermal_expansion):


"\Delta V = \alpha_V\Delta TV_0"

were "\Delta V = V - V_0" is the change in volume, "\Delta T = T - T_0" is the change in temperature, and "\alpha_V = 87\times 10^{-6}\space \degree C^{-1}" is the volumetric expansion coefficient of lead. Thus, obtain:


"V = V_0 + \alpha_V(T-T_0)V_0 = V_0(1 + \alpha_V(T-T_0)) = \dfrac{m}{\rho_0}(1 + \alpha_V(T-T_0))"

The density at "T = 200\degree C" is then:


"\rho = \dfrac{m}{V} = \dfrac{m}{\dfrac{m}{\rho_0}(1 + \alpha_V(T-T_0))} = \dfrac{\rho_0}{1 + \alpha_V(T-T_0)}"

Substituting the values, obtain:


"\rho = \dfrac{11g/cm^3}{1 + 87\times 10^{-6} \space \degree C^{-1}(200\degree C-20\degree C)} \approx 10.8\space g/cm^3"

Answer. "10.8\space g/cm^3".